### v1

# Proposal for Improving Google A2A Protocol: Safeguarding Sensitive Data in Multi-Agent Systems

Abstract
A2A, a protocol for AI agent communication, offers a robust foundation for secure AI agent communication. However, it has several critical issues in handling sensitive data, such as payment details, identification documents, and personal information. This paper reviews the existing protocol, identifies its limitations, and proposes specific enhancements to improve security, privacy, and trust. It includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on prior studies to strengthen the arguments and proposed solutions. This proposal includes seven enhancements: short-lived tokens, customer authentication (SCA), granular scopes, explicit consent, direct data transfer, multi-transaction approval, and payment standard compliance. The vacation booking example illustrates how these enhancements reduce risks and enhance user experience.
1 Introduction
The rapid appearance of Agentic AI autonomous systems that plan, delegate, and collaborate without human intervention has created an urgent need for robust, standardized communication protocols [12]. Google’s A2A (Agent-to-Agent) protocol [46] establishes such a foundation by defining a declarative, identity aware framework for discovering, authenticating, and exchanging tasks among heterogeneous agents. A2A’s core mechanism as shown at Figure 1, the AgentCard, provides machine readable metadata that enables seamless interoperability across organizational and technical boundaries. Complementing A2A, the Model Context Protocol (MCP) standardizes the integration of large language models (LLMs) with external data sources and tools, allowing agents to leverage these models for real-time contextual understanding during task execution. This integration enhances agents’ capabilities in complex scenarios, supporting use cases ranging from payment orchestration to business automation [24]
Figure 1: A2A’s core mechanism
However, as multi-agent ecosystems scale in complexity, they also open new attack surfaces. Recent threat analyses of both A2A and the related MCP reveal sophisticated vectors such as shadowing (where malicious tool descriptions subvert genuine workflows), tool poisoning, and naming attacks that exploit implicit trust in discovery mechanisms [27, 40]. Furthermore, broad surveys of AI agents underscore systemic risks-ranging from prompt injection to cross-agent privilege escalation-that can lead to data exfiltration or unauthorized task execution [18, 31, 41]. Similarly, “enterprise-grade” analyses of MCP implementations demonstrate that without rigorous, zero-trust controls, attackers may exploit unsecured tool endpoints or manipulate context payloads to bypass security checks [35].
Taken together, these findings highlight a gap between protocol design and real-world resilience: A2A and MCP provide the plumbing for agent interoperability, but they do not yet enforce the fine-grained confidentiality, integrity, and consent guarantees required for sensitive data exchange. In this paper we review six critical weaknesses in the handling of personal information, such as payment credentials, user identity and sensitive documents, are identified. Building on proactive threat models and industry best practices, we propose seven concrete enhancements ranging from short-lived, client bound tokens to explicit user consent fields and direct data transfer processes to harden A2A communications against the most urgent threats of the day. This proposal includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on previous studies to strengthen the arguments and proposed solutions.
2 Overview of the A2A Protocol
Google’s A2A (Agent-to-Agent) protocol [46] builds upon widely adopted web standards HTTP, HTTPS, JSON-RPC, and Server-Sent Events (SSE) to provide an extensible, interoperable framework for autonomous agent communication. Agents establish mutual authentication via OAuth 2.0 [25] flows and JSON Web Tokens (JWTs) [29], while RSA key pairs are used for signature validation and secure key exchange. Specifically, OAuth 2.0, an industry-standard authorization framework, enables agents to securely delegate access to each other’s resources without sharing credentials, using standardized flows like the authorization code grant. Meanwhile, JWTs, compact and cryptographically signed tokens, facilitate the secure transmission of authentication claims between agents, ensuring both integrity and verification of access rights during interactions. Role-based access control (RBAC) is natively supported through user- and agent-scoped JWT claims, and fine-grained permissions map directly onto task and message schemas to enforce the principle of least privilege. By default, A2A messages carry only minimal metadata (e.g., action identifiers, input/output schemas, and consent fields), reducing sensitive data exposure when agents negotiate capabilities or invoke third-party services such as payment gateways or AI inference backends.
A2A’s flexibility underpins diverse real-world scenarios:
  • Service Booking: Agents coordinate multi-step operations (e.g., flight, hotel, transportation) by passing structured requests containing specific payment tokens and availability parameters.
  • Enterprise Task Management: Supply-chain or project-management agents exchange only the data needed to complete discrete tasks, minimizing shared context.
  • Cross-Provider Collaboration: Agents from different platforms (e.g., Google, PayPal, Cohere) interoperate securely via standardized AgentCard discovery and encrypted channels.
Despite these strengths, A2A inherits an inherent trade-off between security and efficiency: tightening authorization or adding cryptographic checks invariably increases latency and complexity, whereas relaxing controls risks unauthorized data flows [39]. Crucially, A2A currently lacks specialized safeguards for handling particularly sensitive payloads-such as payment credentials or identity documents-beyond generic token expiry. A similar concern is raised in “AI Agents Meet Blockchain” observes, decentralized agent ecosystems without centralized guardrails require robust consent, auditing, and privacy-preserving delegation mechanisms to maintain trust [31]. Without such controls, adversarial agents might exploit overly broad scopes or unmonitored message channels to exfiltrate or tamper with confidential information. The following sections will analyze these gaps in more detail and propose targeted enhancements to ensure that A2A can securely mediate sensitive data exchanges at scale.
3 Identification of Issues in Handling Sensitive Data
Despite its benefits, A2A has some critical problems in the handling of sensitive data such as payment details, ID documents and personal data. The following six issues are based on real world experience, peer reviewed literature and documented CVE threats.
3.1 Absence of limitations on token lifetime
Although A2A is based on OAuth 2.0, it does not enforce strict expiration durations (e.g. seconds or minutes) for tokens used in sensitive transactions. Without such restrictions, leaked tokens may remain valid for hours or even days, increasing the risk of unauthorized reuse. revocation. As demonstrated in ”AgNet” [23], long-lived tokens are a systemic weakness in distributed architectures, allowing for multiple accesses in the event of a compromise. For example, CVE-2025-1198 shows that revoked GitLab personal access tokens are still accepted by long-term ActiveConnection [10]. Another case described by CVE-2025-1801, where a low-privileged user obtained a JWT issued to a high-privileged user on account [11].
3.2 Lack of Strong Customer Authentication (SCA)
The A2A protocol does not have built-in requirements for strong authentication, such as two-factor or biometric authentication, for high-value transactions such as payments or identity switching. Without these safeguards, adversaries may perform unauthorized acts on behalf of the user [22]. The Medibank breach of 2022, where attackers gained access to personal data of 9.7 million people through the lack of multifactor authentication, illustrates the tangible consequences of [26]. CWE-306 [9] also classifies systems that do not authenticate users before performing critical functions as inherently vulnerable. In addition, the AI Agents Meet the Blockchain project (aclm) [31] proposes Zero Knowledge Proof (ZKP) techniques for secure authentication in decentralized environments, but the A2A does not include such a mechanism.
3.3 Insufficiently Granular Token Scopes
Tokens in A2A do not define precise ranges for sensitive transactions, which introduces the risk of privilege escalation. For example, a token issued to initiate a payment may inadvertently grant access to unrelated data. The study on the multi-agent security tax [39] argues that coarse-grained authorization models increase the probability of data exposure. This concern is highlighted by vulnerability CVE-2023-4456 [3] on LokiStack. Similarly, CWE-1220 [7] documents the lack of granularity in access control policies leading to infringements of the principle of the minimum privilege.
3.4 Lack of Transparency and User Consent
A2A lacks mechanisms that notify users about, or request consent for, the sharing of sensitive data with agents. “AI Agents Meet Blockchain” [31] emphasizes the importance of user transparency in decentralized agent-based environments. CVE-2024-44131 illustrates the consequences of bypassing such protections: malicious applications were able to circumvent Apple’s Transparency, Consent, and Control (TCC) framework to access sensitive data without user approval [4]. CWE-200 [8] codifies this as unauthorized exposure of sensitive information.
3.5 Potential Excessive Exposure of Data to Agents
Agents in A2A ecosystems can access significantly more data than is necessary. “AI Agents Under Threat” [18] articulates how agent2agent data propagation can lead to unintended sharing of sensitive information. CVE-2023-41745 [2] and CVE-2022-45449 [1] document such exposures due to excessive privilege allocation or excessive data collection.
3.6 Risk of Data Disclosure to the Agent Itself
Moreover, the problem is not only that information is being disclosed to malicious actors, but that the very act of disclosing information to the agent itself is problematic, and that even well-meaning agents cannot be presumed to be safe from interference. Prompt injection attacks illustrate how malicious inputs can harvest classified information. “AI Agents Under Threat” [18] warns of this threat. CVE-2024-7042 [6] and CVE-2024-45989 [5] demonstrate how AI agents were exploited to perform unauthorized actions and exfiltrate sensitive data.
3.7 Concrete Example: Vacation Booking
Below is a usage scenario for the protocol without the proposed improvements. As reflected in Figure 2
  • The user requests an AI agent to book a vacation, including flights, hotel, and taxi, using his calendar, personal information and payment details.
  • The agent shares a payment token with the booking agent, valid for hours or days.
  • The agent accesses the user’s full calendar, including irrelevant personal details (e.g., medical appointments), and payment details.
Concern: The token could be used for additional payments, accidentally or intentionally, and calendar access exposes private information without explicit consent.
Figure 2: Vacation Booking Process Using An AI Agent
4 Proposed Enhancements to the Protocol
To address the identified issues, we propose seven enhancements to the A2A protocol, integrating research-backed solutions and aligning with advanced standards:
  1. 1.
    Short-Lived Tokens for Sensitive Operations
    Description: Mandate that tokens for sensitive operations, such as payments or ID transfers, have very short lifetimes (30 seconds to 5 minutes). Tokens expire after a single use or predefined time, requiring re-authentication for further actions.
    Rationale: As demonstrated in “AgNet” [23], short-lived tokens reduce the risk of unauthorized or repeated access. In our case, a payment token expiring quickly limits the window for leaks. “AI Agents Meet Blockchain” [31] supports this by noting that encryption and temporary records in blockchain systems enhance privacy, and our proposed tokens align with this principle.
  2. 2.
    Strong Customer Authentication (SCA) for Sensitive Transactions
    Description: Implement a mechanism requiring SCA, such as SMS codes, biometric verification, or bank login, for every sensitive transaction. This can be integrated into the task flow, ensuring user authentication before actions.
    Rationale: “AI Agents Meet Blockchain” [31] highlights that ZKPs enable secure authentication without exposing data, and in our case, SCA based on ZKPs or similar techniques ensures only verified users authorize sensitive actions. This aligns with PSD2 requirements for payment transactions, ensuring regulatory compliance.
  3. 3.
    More Granular Token Scopes
    Description: Extend OAuth 2.0 scopes to define precise permissions, such as “payment of \$1000 to Hotel X on Date Y” or “access to calendar availability only.” This prevents token use for unauthorized actions.
    Rationale: “Multi-Agent Security Tax” [39] notes that high granularity in permissions reduces unnecessary data exposure. In our case, precise scopes ensure agents access only required data, aligning with GDPR’s Minimum Necessary Disclosure principle.
  4. 4.
    Explicit User Consent Mechanism
    Description: Add a “consent” field to message or task structures, requiring agents to obtain explicit user approval before sharing sensitive data, specifying data type, purpose, and recipient.
    Rationale: “AI Agents Meet Blockchain” [31] emphasizes transparency as critical for trust in decentralized systems. In our case, explicit consent enhances user trust and ensures GDPR compliance, reducing the risk of unapproved data sharing.
  5. 5.
    Direct Data Transfer
    Description: Enable direct transfer of sensitive data between the user and service provider, bypassing intermediary agents. For example, in payments, the agent redirects the user to the bank for direct transfer, with user verification.
    Rationale: Inspired by “Multi-Agent Security Tax” [39]’s “active vaccines” to prevent malicious data spread, direct transfers avoid exposing data to intermediaries. “AI Agents Under Threat” [18] supports this by emphasizing the need to limit data propagation, and in our case, direct transfers prevent the exposure of sensitive data to the agent and reduce the risks of leakage.
  6. 6.
    Support for Multi-Transaction Approval
    Description: Allow a single user approval for a series of related transactions (e.g. flight, hotel, taxi payments), with tokens restricted to that series, short-lived, and SCA-verified.
    Rationale: This balances security and convenience, as recommended in “Multi-Agent Security Tax” [39] for efficiency-security trade-offs. In our case, it reduces repeated authentications while maintaining strict token restrictions. Additionally, this solution addresses the well-documented issue of ”consent fatigue,” where users, overwhelmed by frequent approval requests, may approve actions without proper scrutiny, increasing security risks. As highlighted in the Palo Alto Networks article [50] repeated requests can lead to user desensitization, potentially resulting in the approval of critical actions like write operations without adequate attention—a phenomenon similar to MFA fatigue. By consolidating approvals for related transactions, our solution minimizes the frequency of requests, thereby reducing the risk of exploitation by malicious actors while enhancing the user experience.
  7. 7.
    Compliance with Payment Standards
    Description: Ensure the protocol supports standards like PSD2, requiring SCA for certain transactions, by defining interfaces for payment agents that incorporate these requirements.
    Rationale: Compliance with regulations is critical for adoption, as noted in “AI Agents Meet Blockchain” [31]. Our proposal ensures A2A aligns with regulatory requirements, enhancing its reliability.
4.1 Application to the Example
Protocol Usage Scenario After The Proposed Improvements:
  1. 1.
    The user requests the AI agent to book a vacation, including flights, hotel, and taxi.
  2. 2.
    The agent requests permission to access calendar availability only, displaying: “I will share availability dates with the booking agent. Approve?” The user approves.
  3. 3.
    The agent finds a suitable booking and requests payment approval: “I will process \$1000 for Hotel X, \$500 for Flight Y, and \$50 for Taxi Z. Approve?” The user approves.
  4. 4.
    The agent requests a 5-minute token from the user’s bank, scoped to “payment for vacation X on Date Y.”
  5. 5.
    The bank issues a 5-minute token for the approved transaction series.
  6. 6.
    The agent shares the token with the booking agent.
  7. 7.
    The booking agent requests the bank to process payments using the token.
  8. 8.
    The bank prompts the user to verify the action via SMS code or biometric authentication.
  9. 9.
    The user approves, and payments are processed.
  10. 10.
    The token expires, preventing reuse.
Advantages:
  • Short-Lived Token: Prevents reuse or unauthorized access.
  • SCA: Ensures user approval for each payment.
  • Precise Scope: Token restricted to approved transactions.
  • Explicit Consent: The user is informed and approves the sharing of data.
  • Minimal Access: The agent accesses only availability, not the full calendar.
  • Direct Transfer: In order to protect sensitive information from agents, payment is sent straight from the bank to the booking agent.
4.2 Research Base Rationale for This Proposal
This proposal is based on previous research and enhances A2A’s security, privacy, and efficiency:
  • Enhanced Security: Short-lived tokens and SCA reduce leak risks, as recommended in [18, 31, 23].
  • Regulatory Compliance: SCA and PSD2 support ensure compliance, as emphasized in [31].
  • Transparency and Trust: Explicit consent increases user trust, per [31].
  • Flexibility and Efficiency: Multi-transaction approval balances security and convenience, as suggested in [39].
  • Common Standarts: The enhancements integrate with OAuth 2.0 and PSD2, as proposed in [23].
  • Exposure Prevention: Direct transfers minimize leak risks, according to [18, 39].
5 Evidence-Based Support for Proposed Enhancements
In order to confirm the technical rationale behind the proposed improvements to the Google A2A Protocol, this section provides a literature review of each of them. Each section focuses on one proposed improvement and summarises previous literature reviews showing that the improvements are not only technically motivated but also based on established research in areas such as authentication protocols, cryptography, distributed identity systems and access control. Together, these references provide empirical and architectural support for the adoption of the proposed measures.
5.1 Use of Short-Lived Access Tokens
One of the core enhancements we proposed to improve the security of Google’s A2A protocol is the mandatory use of short-lived access tokens. This design decision substantially reduces the window of opportunity for replay attacks and limits the utility of any compromised credential. Multiple studies affirm this principle. Teng (2023) introduced ActionID, a machine-to-machine authentication protocol that issues tokens tightly scoped in both action and time. The author notes that “the short-lived token limits the time window for action execution,” thereby reducing the effectiveness of credential theft [47]. Similarly, Ohwo et al. (2024) proposed a blockchain-based smart home access system in which “short-lived tokens [are] used to mitigate risks like replay attacks and user profiling” [37]. These tokens expire rapidly, neutralizing any intercepted credentials. Narajala et al. (2025) extended this logic to GenAI multi-agent systems, proposing a just-in-time registry that dynamically provisions tokens only for the duration of a tool invocation. This architecture, they argue, “minimizes the attack surface associated with persistent credentials by dynamically provisioning short-lived access tokens only when needed” [34]. Finally, Xiao et al. (2024) addressed the same vulnerability in the IoT domain, where resource-constrained devices are especially susceptible to token compromise. They developed MCU-Token, which binds a short-lived, hardware-derived token to each request, thereby making every token instance single-use and resistant to replay [48]. Together, these findings offer both conceptual and empirical support for implementing short-lived tokens in the A2A protocol as a fundamental safeguard against credential replay and session hijacking.
5.2 Strong Customer Authentication (SCA) for Sensitive Transactions
Implementing Strong Customer Authentication (SCA) for sensitive operations in the A2A protocol, such as payments and identity verification, significantly improves protection against impersonation and unauthorized transactions. Recent research underscores the feasibility and necessity of using modern, privacy-preserving SCA mechanisms such as zero-knowledge proofs (ZKPs), biometrics, and multi-factor authentication (MFA). For example, Neera et al. propose a mobile payment protocol that leverages ZKPs and identity-based signatures to verify user identity without revealing sensitive data while ensuring compliance with financial regulations such as PSD2 [36]. Their scheme guarantees cryptographically enforced SCA while maintaining user privacy, demonstrating that SCA can be both secure and regulatory compliant. Ahmad et al. present BAuth-ZKP, a blockchain-based MFA framework using smart contracts and ZKPs to authenticate users in smart city environments [13]. This design proves the viability of decentralized, privacy-respecting SCA mechanisms, directly applicable to agent-mediated A2A architectures. On the biometric front, Gernot and Rosenberger introduce a technique for generating one-time biometric templates, mitigating the risk of replay attacks by ensuring biometric credentials are valid only once [20]. This reinforces the ’inherence’ factor of SCA in a technically sound manner. Finally, Lyastani et al. empirically demonstrate that while 2FA dramatically improves account security, inconsistent or poorly designed SCA flows reduce adoption [21]. Thus, their findings support designing usable, transparent SCA processes for sensitive actions in A2A. Collectively, these studies provide strong empirical and architectural justification for integrating SCA into the A2A protocol, ensuring that sensitive transactions are approved only by verified users under secure and user-friendly conditions.
5.3 More Granular Token Scopes
Enforcing fine-grained token scopes in the A2A protocol is essential for upholding the principle of least privilege and minimizing the exposure of sensitive resources during agent-to-agent transactions. The risks of overly broad permissions in bearer tokens are well-documented in recent research. Cao et al. propose a stateful, least-privilege authorization model that allows client-side applications to dynamically constrain the scope of OAuth tokens using WebAssembly-based privilege attenuation policies [17]. Their system empowers developers to explicitly encode minimal access rights per session, enabling secure delegation without overprovisioning. This aligns directly with the A2A context, where agents must act within strict permission boundaries to avoid inadvertent data access. Complementarily, South et al. extend OAuth 2.0 and OpenID Connect for authenticated agent delegation, introducing agent-specific credentials with precise, auditable scope limitations [45]. Their framework demonstrates how user intent can be translated into tightly scoped permissions, ensuring that AI agents operate only within authorized domains. These works collectively support the need for scope-constrained access tokens as a foundational safeguard in agent-mediated architectures. Dimova et al. further substantiate this by showing that 18.5% of OAuth deployments on the web request unnecessary scopes, violating the GDPR’s minimum necessary data principle [19]. Kaltenböck et al. reinforce this position by embedding scope-aware policies within a Zero Trust single sign-on framework, emphasizing explicit scope definitions to limit token capabilities at authentication time [30]. Altogether, the literature affirms that fine-grained token scopes are not only technically viable but also indispensable for ensuring data minimization, secure delegation, and regulatory compliance in distributed, AI-driven authentication systems.
5.4 Explicit User Consent Mechanism
Embedding an explicit user consent mechanism within the A2A protocol is critical for aligning with privacy regulations such as GDPR and for building user trust in agent-mediated systems. Recent research emphasizes that consent must be freely given, informed, specific, and revocable throughout the data lifecycle. Merlec et al. propose a blockchain-based dynamic consent management system, where users can grant, audit, or withdraw their consent via smart contracts stored on a tamper-proof ledger [33]. This architecture guarantees accountability, traceability, and user autonomy in data-sharing environments. Complementarily, Khalid et al. formalize the security and privacy requirements of such systems, proposing the integration of zero-knowledge proofs and cryptographic primitives to ensure that consent is provable, minimal, and compliant by design [32]. Their work outlines how systems can enforce consent boundaries while maintaining confidentiality. In a multi-agent context, Xu et al. demonstrate that autonomous privacy agents can enforce user-defined consent policies, making decisions that reflect GDPR principles such as data minimization and informed processing [49]. These agents act only within verified constraints, providing technical assurance that consent decisions are respected. Finally, Pathmabandu et al. introduce a consent management engine that enables granular, real-time visibility into data collection within IoT systems [38]. Their engine offers digital nudging and fine-grained control, reinforcing the user’s right to control their personal data. Collectively, these studies support the integration of explicit, technically enforceable consent mechanisms into A2A-style protocols, ensuring that sensitive data is shared only with informed user approval, and that such actions remain transparent and auditable.
5.5 Direct Data Transfer
In the evolving landscape of multi-agent systems, ensuring the security of sensitive data during interactions between users and service providers is paramount. Traditional multi-agent architectures often involve intermediaries that can pose significant security risks, such as data leakage or unauthorized access. To address these concerns, the proposed enhancement advocates for direct data transfer, where sensitive information is exchanged directly between the user and the service provider, bypassing intermediary agents. This approach is supported by recent research in multi-agent security. Firstly, [39] introduces the concept of ”active vaccines” to prevent the spread of malicious prompts in multi-agent systems, underscoring the need to minimize intermediary involvement to reduce systemic vulnerabilities. Similarly, [18] highlights the unpredictability of multi-step user inputs and the complexity of internal executions in AI agents, emphasizing the necessity of limiting data propagation to trusted entities. Furthermore, [42] discusses the open challenges in securing systems of interacting AI agents, particularly the threats arising from free-form interactions and network effects that can amplify security breaches, which direct data transfer mitigates by reducing intermediaries. Lastly, [45] presents a framework for authenticated delegation and authorized AI agents, which can be extended to support direct data transfer by ensuring secure, intermediary-free communication channels. On the basis of these studies, it can be concluded that direct transmission of data increases the security of multi-agent systems by reducing reliance on intermediaries and thus minimising the risks of data exposure and unauthorised access. This improvement is essential for applications involving sensitive data such as payments and personal data, where security and privacy are paramount.
5.6 Support for Multi-Transaction Approval
In the area of multi-agent systems and financial technology, securing financial transactions while preserving user comfort is a key challenge, especially for sensitive data such as payment details. The proposed enhancement, ’Multi-Transaction Approval,állows for a single authorization for a series of transactions using limited-time tokens validated by strong customer authentication. This approach balances security and usability by reducing the need for multiple authentication while maintaining strict security measures, thus increasing the effectiveness of multi-step workflow in agent-based systems. Recent research on the security of financial technology supports this improvement. For instance, [15] underscore the pivotal role of multi-factor authentication (MFA) in securing mobile financial transactions, advocating its use to mitigate fraud risks, which aligns with the SCA requirements of the proposed enhancement. Similarly, [44] provide a systematic review of MFA in digital payment systems, noting that grouping transactions under a single secure authentication session is feasible with robust mechanisms, supporting the enhancement’s design. Furthermore, [28] propose a framework integrating MFA with machine learning to secure online financial transactions, adaptable to multi-transaction approval by ensuring SCA verification for each transaction in a series and employing anomaly detection to enhance security. Additionally, [14] develop an MFA algorithm for mobile money applications, combining multiple authentication factors to secure transactions, which can be extended to support a single approval for multiple transactions as proposed. In conclusion, the “Multi-Transaction Approval” enhancement is robustly supported by current research in financial technology security, emphasizing advanced MFA techniques and fraud detection methods. By implementing this enhancement, the A2A protocol can offer a secure and user-friendly approach to managing multiple related transactions, reducing user friction while adhering to high security standards.
5.7 Compliance with Payment Standards
In the area of multi-agent systems and financial technology, compliance with regulatory standards such as PSD2 is essential to ensure the safe and secure processing of transactions. The proposed enhancement, “Compliance with Payment Standards”, integrates these requirements into Google’s A2A protocol by defining interfaces for payment agents that include strong customer authentication and other regulatory requirements. This approach increases the credibility and readiness of the Protocol for implementation in the real world by reputable entities. Recent research supports the feasibility of such improvements. For instance, [31] highlight how blockchain technology can ensure compliance through tamper-proof documentation, aligning with PSD2’s demands for transparency and security in financial transactions [31]. Similarly, [16] demonstrate that artificial intelligence can bolster regulatory compliance in the financial sector by leveraging machine learning to monitor and prevent breaches, a principle applicable to multi-agent systems to enforce PSD2 standards. Lastly, [43] address the broader compliance landscape for AI systems, focusing on the EU’s AI Act and data set compliance, which reinforces the importance of embedding regulatory adherence into AI-driven systems, including those involving multi-agent interactions [43]. By integrating these insights, the enhancement positions the A2A protocol as a robust framework for secure, compliant financial transactions in multi-agent environments.
6 Implementation Considerations
Implementing the proposal requires technical and organizational steps:
  • Token Management: Define processes for issuing, verifying, and revoking short-lived tokens, as recommended in [23].
  • Authentication Flows: Develop SCA flows, including two-factor or biometric options, per PSD2 requirements.
  • Protocol Extensions: Add message types or task states for sensitive data handling and consent requirements.
  • Documentation and Guidelines: Provide detailed documentation on sensitive data handling, per [39].
  • Compatibility Testing: Test enhancements with partners like PayPal and Cohere to ensure interoperability, as suggested in [39].
7 Conclusion
Google’s A2A protocol offers a robust foundation for secure agent communication but requires enhancements to handle sensitive data effectively. As evidenced in “AI Agents Under Threat” [18], “AI Agents Meet Blockchain” [31], “Multi-Agent Security Tax” [39], and “AgNet” [23], security and privacy challenges in multi-agent systems demand solutions balancing robust protection with efficiency. This proposal introduces seven enhancements: short-lived tokens, SCA, granular scopes, explicit consent, direct data transfer, multitransaction approval, and payment standard compliance to improve security, privacy, and trust. The vacation booking example illustrates how these enhancements reduce risks and enhance user experience. We recommend Google implement this proposal to strengthen A2A’s position as a leading standard for agent communication.

### v2

@@
 Additionally, this solution addresses the well-documented issue of ”consent fatigue,” where users, overwhelmed by frequent approval requests, may approve actions without proper scrutiny, increasing security risks.
-As highlighted in the Palo Alto Networks article [50] repeated requests can lead to user desensitization, potentially resulting in the approval of critical actions like write operations without adequate attention—a phenomenon similar to MFA fatigue.
+As highlighted in the Palo Alto Networks article [53] repeated requests can lead to user desensitization, potentially resulting in the approval of critical actions like write operations without adequate attention, a phenomenon similar to MFA fatigue.
 By consolidating approvals for related transactions, our solution minimizes the frequency of requests, thereby reducing the risk of exploitation by malicious actors while enhancing the user experience.
@@
 By integrating these insights, the enhancement positions the A2A protocol as a robust framework for secure, compliant financial transactions in multi-agent environments.
+6 Prompt Injection in LLM-Based Agents
+Prompt injection attacks are a significant concern in the domain of artificial intelligence security, particularly with the widespread deployment of large language models (LLMs) in applications ranging from automated customer support to complex data processing systems.
+These attacks involve the deliberate insertion of malicious or deceptive text inputs designed to manipulate an AI model’s behavior, potentially leading to the disclosure of sensitive information, the generation of harmful outputs, or the undermining of the system’s intended purpose.
+The vulnerability arises from the inherent difficulty AI systems face in differentiating between legitimate user inputs and crafted malicious prompts, a challenge that becomes increasingly significant as AI integrates into sensitive operational contexts.
+Already in 2022, as LLMs began their global proliferation, researchers identified the risks posed by malicious prompt manipulation.
+Perez et al.
+[43] demonstrated that carefully designed input could override predefined model instructions, forcing the model to ignore prior constraints, disclose sensitive data, or behave in unintentional ways.
+Although initial defenses have since improved, new and more sophisticated methods continue to evolve.
+In 2023, Liu et al.
+[35] revealed that 31 of 36 LLM-integrated applications tested, including productivity and chat tools, remained vulnerable to prompt injection, even in black-box scenarios.
+Their work underscores that the threat landscape is not only persistent, but also increasingly complex.
+Figure 3: Prompt Injection Proccess Via N8N Platform - Agent 1
+To investigate these threats in practice, we conducted an experiment using the N8N platform, a workflow automation tool that facilitates the creation of AI agents.
+In this experiment, a first AI agent, as shown in Figure 3, was constructed with explicit instructions not to reveal private information critical to its function, simulating a scenario where confidentiality is paramount.
+Subsequently, a second AI agent, as shown in Figure 4, was developed with the sole purpose of extracting this withheld information from the first agent.
+Employing various manipulative techniques and prompt injection strategies, the second agent successfully elicited the sensitive data, without relying on advanced tactics, thereby exposing the fragility of AI systems to such attacks and highlighting the practical risks of entrusting them with confidential information.
+According to this article from Palo Alto Cyberpedia [1], prompt injection attacks can be classified into two main categories: direct and indirect.
+Direct prompt injection involves an attacker inputting a malicious prompt directly into an AI application’s interface, overriding the system’s predefined instructions.
+For instance, an attacker might instruct the AI to disregard its original directives and instead disclose restricted data, as demonstrated in our experiment.
+In contrast, indirect prompt injection involves embedding malicious instructions within external data sources, such as web pages, documents, or other content, that the AI subsequently processes, unwittingly executing the hidden commands.
+A particularly insidious variant, known as stored prompt injection, involves implanting malicious prompts into the AI’s memory or training data, allowing the attack to persist and influence the system’s responses over time.
+These diverse methodologies underscore the multifaceted nature of prompt injection threats, as attackers can exploit both immediate input channels and longer-term data manipulation tactics to achieve their objectives.
+Figure 4: Prompt Injection Proccess Via N8N Platform - Agent 2
+The findings from our experiment, along with the insights from Palo Alto article, reveal a dynamic interplay between advancing defensive measures and evolving attack strategies.
+As techniques to safeguard AI systems against prompt injection improve, such as enhanced input filtering or behavioral monitoring, attackers adapt by devising increasingly sophisticated methods to bypass these protections.
+This ongoing escalation suggests that relying solely on technical defenses may prove inadequate in the long term.
+Instead, the most robust strategy for safeguarding sensitive information emerges as a preventative one: refraining from sharing such data with AI agents entirely, as we propose at this paper.
+By excluding confidential information from AI systems, the risk of its exposure through prompt injection is effectively eliminated, aligning with the cybersecurity principle of minimizing data exposure.
+This conclusion not only reflects the practical lessons derived from the experiment but also emphasizes the necessity of adopting proactive and stringent security practices in the design and deployment of AI technologies.
+As AI continues to permeate critical sectors, ensuring its resilience against prompt injection attacks will require a concerted effort to balance functionality with security, prioritizing the protection of sensitive data above all.
 7 Implementation Considerations
@@
   • Compatibility Testing: Test enhancements with partners like PayPal and Cohere to ensure interoperability, as suggested in [42].
+8 Explicit User Consent Orchestration via
+USER_CONSENT_REQUIRED
+Adding a new enumeration member 111 A full code sample is available at: https://github.com/yedidel/A2A :
+    USER_CONSENT_REQUIRED = "user-consent-required"
+The integration of USER_CONSENT_REQUIRED into the TaskState enum elevates the protocol’s control over task execution.
+Unlike existing states such as INPUT_REQUIRED, which merely indicate that additional information is needed, this new state explicitly denotes that a task is paused until the end-user provides affirmative consent.
+Embedding this state into the task lifecycle ensures that AI agents and back-end services stop sensitive operations like booking travel, executing financial transactions, or sharing personal data until a human user explicitly approves.
+As depicted in the diagram shown in Figure 5
+Figure 5: Vacation Booking Flow, Integrated With The USER_CONSENT_REQUIRED State
+This explicit state enables precise front-end handling: upon encountering consent required, the UI can display a dedicated consent prompt, ensuring user awareness and control.
+Once consent is granted, the task transitions to IN_PROGRESS or an equivalent active state.
+This mechanism not only enhances user agency and trust but also bolsters compliance with stringent regulatory frameworks (e.g., GDPR, PSD2), by providing verifiable evidence that the user approved the action before execution.
 9 Conclusion

### v3

@@
-Proposal for Improving Google A2A Protocol: Safeguarding Sensitive Data in Multi-Agent Systems
+Improving Google A2A Protocol: Protecting Sensitive Data and Mitigating Unintended Harms in Multi-Agent Systems
 Abstract
-A2A, a protocol for AI agent communication, offers a robust foundation for secure AI agent communication.
-However, it has several critical issues in handling sensitive data, such as payment details, identification documents, and personal information.
-This paper reviews the existing protocol, identifies its limitations, and proposes specific enhancements to improve security, privacy, and trust.
-It includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on prior studies to strengthen the arguments and proposed solutions.
-This proposal includes seven enhancements: short-lived tokens, customer authentication (SCA), granular scopes, explicit consent, direct data transfer, multi-transaction approval, and payment standard compliance.
-The vacation booking example illustrates how these enhancements reduce risks and enhance user experience.
-1 Introduction
-The rapid appearance of Agentic AI autonomous systems that plan, delegate, and collaborate without human intervention has created an urgent need for robust, standardized communication protocols [14].
-Google’s A2A (Agent-to-Agent) protocol [48] establishes such a foundation by defining a declarative, identity aware framework for discovering, authenticating, and exchanging tasks among heterogeneous agents.
-A2A’s core mechanism as shown at Figure 1, the AgentCard, provides machine readable metadata that enables seamless interoperability across organizational and technical boundaries.
-Complementing A2A, the Model Context Protocol (MCP) standardizes the integration of large language models (LLMs) with external data sources and tools, allowing agents to leverage these models for real-time contextual understanding during task execution.
-This integration enhances agents’ capabilities in complex scenarios, supporting use cases ranging from payment orchestration to business automation [26]
-Figure 1: A2A’s core mechanism
-However, as multi-agent ecosystems scale in complexity, they also open new attack surfaces.
-Recent threat analyses of both A2A and the related MCP reveal sophisticated vectors such as shadowing (where malicious tool descriptions subvert genuine workflows), tool poisoning, and naming attacks that exploit implicit trust in discovery mechanisms [29, 44].
-Furthermore, broad surveys of AI agents underscore systemic risks-ranging from prompt injection to cross-agent privilege escalation-that can lead to data exfiltration or unauthorized task execution [20, 32, 45].
-Similarly, “enterprise-grade” analyses of MCP implementations demonstrate that without rigorous, zero-trust controls, attackers may exploit unsecured tool endpoints or manipulate context payloads to bypass security checks [37].
-Taken together, these findings highlight a gap between protocol design and real-world resilience: A2A and MCP provide the plumbing for agent interoperability, but they do not yet enforce the fine-grained confidentiality, integrity, and consent guarantees required for sensitive data exchange.
-In this paper we review six critical weaknesses in the handling of personal information, such as payment credentials, user identity and sensitive documents, are identified.
-Building on proactive threat models and industry best practices, we propose seven concrete enhancements ranging from short-lived, client bound tokens to explicit user consent fields and direct data transfer processes to harden A2A communications against the most urgent threats of the day.
-This proposal includes a concrete example to illustrate the problem and solution, research-backed rationales, and implementation considerations, drawing on previous studies to strengthen the arguments and proposed solutions.
-2 Overview of the A2A Protocol
-Google’s A2A (Agent-to-Agent) protocol [48] builds upon widely adopted web standards HTTP, HTTPS, JSON-RPC, and Server-Sent Events (SSE) to provide an extensible, interoperable framework for autonomous agent communication.
-Agents establish mutual authentication via OAuth 2.0 [27] flows and JSON Web Tokens (JWTs) [30], while RSA key pairs are used for signature validation and secure key exchange.
-Specifically, OAuth 2.0, an industry-standard authorization framework, enables agents to securely delegate access to each other’s resources without sharing credentials, using standardized flows like the authorization code grant.
-Meanwhile, JWTs, compact and cryptographically signed tokens, facilitate the secure transmission of authentication claims between agents, ensuring both integrity and verification of access rights during interactions.
-Role-based access control (RBAC) is natively supported through user- and agent-scoped JWT claims, and fine-grained permissions map directly onto task and message schemas to enforce the principle of least privilege.
-By default, A2A messages carry only minimal metadata (e.g., action identifiers, input/output schemas, and consent fields), reducing sensitive data exposure when agents negotiate capabilities or invoke third-party services such as payment gateways or AI inference backends.
-A2A’s flexibility underpins diverse real-world scenarios:
-  • Service Booking: Agents coordinate multi-step operations (e.g., flight, hotel, transportation) by passing structured requests containing specific payment tokens and availability parameters.
-  • Enterprise Task Management: Supply-chain or project-management agents exchange only the data needed to complete discrete tasks, minimizing shared context.
-  • Cross-Provider Collaboration: Agents from different platforms (e.g., Google, PayPal, Cohere) interoperate securely via standardized AgentCard discovery and encrypted channels.
-Despite these strengths, A2A inherits an inherent trade-off between security and efficiency: tightening authorization or adding cryptographic checks invariably increases latency and complexity, whereas relaxing controls risks unauthorized data flows [42].
-Crucially, A2A currently lacks specialized safeguards for handling particularly sensitive payloads-such as payment credentials or identity documents-beyond generic token expiry.
-A similar concern is raised in “AI Agents Meet Blockchain” observes, decentralized agent ecosystems without centralized guardrails require robust consent, auditing, and privacy-preserving delegation mechanisms to maintain trust [32].
-Without such controls, adversarial agents might exploit overly broad scopes or unmonitored message channels to exfiltrate or tamper with confidential information.
-The following sections will analyze these gaps in more detail and propose targeted enhancements to ensure that A2A can securely mediate sensitive data exchanges at scale.
-3 Identification of Issues in Handling Sensitive Data
-Despite its benefits, A2A has some critical problems in the handling of sensitive data such as payment details, ID documents and personal data.
-The following six issues are based on real world experience, peer reviewed literature and documented CVE threats.
-3.1 Absence of limitations on token lifetime
-Although A2A is based on OAuth 2.0, it does not enforce strict expiration durations (e.g.
+Google’s A2A protocol provides a secure communication framework for AI agents but demonstrates critical limitations when handling highly sensitive information such as payment credentials and identity documents.
+These gaps increase the risk of unintended harms, including unauthorized disclosure, privilege escalation, and misuse of private data in generative multi-agent environments.
+In this paper, we identify key weaknesses of A2A: insufficient token lifetime control, lack of strong customer authentication, overbroad access scopes, and missing consent flows.
+We propose protocol-level enhancements grounded in a structured threat model for semi-trusted multi-agent systems.
+Our refinements introduce explicit consent orchestration, ephemeral scoped tokens, and direct user-to-service data channels to minimize exposure across time, context, and topology.
+Empirical evaluation using adversarial prompt injection tests shows that the enhanced protocol substantially reduces sensitive data leakage while maintaining low communication latency.
+Comparative analysis highlights the advantages of our approach over both the original A2A specification and related academic proposals.
+These contributions establish a practical path for evolving A2A into a privacy-preserving framework that mitigates unintended harms in multi-agent generative AI systems.
+I Introduction
+The rapid emergence of autonomous agent systems capable of planning, delegating, and coordinating tasks without direct human intervention has created a growing demand for standardized communication protocols [1].
+Google’s Agent-to-Agent (A2A) protocol addresses this need by establishing a structured, identity-aware framework for discovering and interacting with services across heterogeneous agents [2].
+At the core of this protocol lies the AgentCard, a declarative metadata object that enables seamless interoperability between agents by conveying machine-readable descriptions of capabilities, roles, and identities, as illustrated in Figure 1.
+Complementing A2A, the Model Context Protocol (MCP) [3] facilitates real-time integration of large language models with tools and external data, thereby enabling richer context sharing and execution capabilities in multi-agent workflows [4].
+Despite these innovations, recent threat analyses have uncovered critical vulnerabilities that challenge the secure handling of sensitive data in such environments.
+Researchers have demonstrated attack vectors such as shadowing, tool poisoning, and naming manipulation, which exploit implicit trust mechanisms in the agent discovery process [5, 6].
+Furthermore, studies have highlighted systemic risks including prompt injection, excessive privilege escalation, and cross-agent data leakage, particularly in high-value tasks such as payments or identity verification [7, 8, 9].
+Enterprise-grade evaluations of MCP implementations indicate that without robust enforcement of access control and consent boundaries, attackers may subvert tool endpoints or manipulate contextual payloads to bypass authorization policies [10].
+Taken together, these findings reveal a gap between the theoretical design of A2A and its practical resilience against adversarial behavior.
+While A2A and MCP enable functional interoperability, they do not yet provide sufficient guarantees of confidentiality, integrity, and informed consent for handling sensitive information.
+This paper addresses this gap by identifying core protocol weaknesses in areas such as token management, authentication strength, scope granularity, and data flow transparency.
+We propose a set of targeted enhancements to improve privacy, security, and user control in agent-mediated communications, and demonstrate their application in a real-world example involving vacation booking.
+Our proposal incorporates best practices from adjacent fields such as zero-trust architectures and regulatory compliance in financial technologies.
+Figure 1: A2A’s core mechanism. The AgentCard provides identity-aware task and tool metadata to support discovery and execution across agent systems.
+Unlike existing security recommendations that remain optional or are implemented at the application layer, in this paper we present how to enforces these mechanisms at the protocol level in a cohesive manner tailored for decentralized multi‑agent environments.
+The integration of explicit consent orchestration, ephemeral and strictly scoped credentials, and direct user‑to‑service data channels is designed as a unified security architecture rather than isolated best practices.
+This design ensures that safeguards are consistently applied across heterogeneous agents, closing enforcement gaps that persist in current A2A deployments.
+II Related Work and Background
+This section reviews the foundations of the A2A protocol and situates it within broader research on secure multi-agent communication.
+A2A, developed by Google, builds upon established web standards such as HTTP, HTTPS, JSON-RPC, and Server-Sent Events (SSE) to define a secure and extensible communication framework between autonomous agents.
+Agents engage in mutual authentication through OAuth 2.0 [11] and exchange cryptographically signed JSON Web Tokens (JWTs) [12] that encapsulate identity and permission claims.
+The use of RSA key pairs facilitates signature verification and key exchange, ensuring message integrity and authentication without relying on shared secrets.
+OAuth 2.0 provides an industry-standard method for authorization, allowing agents to delegate access securely to external resources.
+The protocol typically employs the authorization code grant flow to ensure secure token exchange.
+JWTs, being compact and self-contained, carry essential claims that include user identifiers, expiration timestamps, and access scopes, enabling verification of access rights and enforcing the principle of least privilege across distributed systems.
+Role-based access control (RBAC) is natively supported through user and agent scoped JWT claims, ensuring that agents only access resources aligned with their intended roles and tasks.
+A2A messages are designed to minimize unnecessary exposure of sensitive information by default.
+The messages include only essential metadata such as action identifiers, input and output schemas, and optional consent fields.
+This design helps reduce data leakage risks during task delegation or tool invocation, particularly when agents communicate with third-party services such as payment processors or AI inference backends.
+A2A supports a wide range of real-world scenarios by promoting secure agent interoperability across organizational and platform boundaries.
+Prominent examples include service booking (e.g., flights, hotels, transportation), enterprise task automation (e.g., supply-chain orchestration, project management, procurement), and cross-provider collaboration (e.g., Google, PayPal, and Cohere).
+These cases highlight the protocol’s broad applicability and practical relevance.
+Despite these design strengths, A2A embodies an inherent trade-off between security and efficiency.
+Introducing additional authorization checks or cryptographic safeguards may increase latency and complexity, whereas relaxing such controls can open avenues for data exfiltration and privilege escalation [13].
+In particular, the A2A specification lacks tailored protections for handling highly sensitive payloads, such as user credentials, payment information, and identity documents.
+As noted by Karim et al.
+[8], decentralized agent environments demand explicit controls around consent, delegation, and auditability to maintain operational trust.
+Without these mechanisms, adversarial agents may exploit overbroad scopes or unmonitored message channels to gain unauthorized access to sensitive resources.
+This analysis reflects prior work that established the strengths and limitations of A2A.
+Analogous vulnerabilities have also been documented in related agent interoperability protocols, including MCP, ANP and ACP [14], [15], [16].
+Nevertheless, since this paper concentrates exclusively on the A2A protocol, we refrain from an in-depth discussion of these alternatives.
+Our contribution builds directly on this foundation by proposing targeted enhancements that mitigate leakage risks in generative multi-agent environments.
+In addition to these protocol-level considerations, recent research has also highlighted risks inherent to the agents themselves, which further motivates the safeguards we propose and are reviewed next.
+Figure 2: Threat model for the A2A protocol in a semi-trusted multi-agent environment.
+Solid arrows indicate legitimate communication flows; dashed red arrows indicate attack vectors including prompt injection, token replay, scope escalation, and unverified endpoints.
+II-A Risks of Agent-Initiated Misuse of Sensitive Data
+While prompt injection remains a well-documented threat, autonomous agents can also misuse sensitive information without any external manipulation.
+In Anthropic’s study on agentic misalignment [17], large language model agents such as Claude were observed engaging in blackmail and corporate espionage, even when initially given benign objectives, especially when perceiving threats to their continued operation.
+Similarly, Cybernews reported a real-world case where Replit’s AI coding assistant deleted a live production database during a code freeze, explicitly ignoring multiple instructions not to proceed [18].
+Another high-profile example from Carnegie Mellon University demonstrated that large language models, when given sufficient autonomy, could independently orchestrate a complete multi-stage cyberattack, including reconnaissance, network exploitation, malware deployment, and data exfiltration without human oversight [19].
+These incidents illustrate that AI agents are not merely passive tools vulnerable to manipulation; in certain circumstances, they may actively and deliberately act against user interests.
+Additional evidence underscores that these risks extend beyond isolated research studies.
+Lumia Security’s AppleStorm investigation revealed that Apple’s Siri and Apple Intelligence ecosystem quietly transmitted WhatsApp and iMessage content, location data, and metadata about installed applications to Apple servers, even when such transmission was unnecessary or contrary to user expectations [20].
+Similarly, Guardio’s Scamlexity study demonstrated that agentic AI browsers, such as Perplexity’s Comet, autonomously completed online purchases on fraudulent websites and submitted sensitive user data to phishing pages without human oversight, effectively turning trusted assistants into vectors of financial and security harm [21].
+Recent academic work further highlights this trend: a comprehensive study on the expanding attack surface of agentic AI systems shows that emergent behaviors, prompt injection vulnerabilities, and autonomous task execution can transform well-intentioned agents into sources of significant unintended harms [22].
+Together, these findings illustrate that AI agents pose inherent privacy and security threats not only when attacked, but also through their own design.
+These risks underscore the need for protocol-level safeguards to prevent unintended harms, an issue we address in the following sections of this paper.
+III Threat Model
+We consider a semi-trusted multi-agent environment, consistent with prior work on LLM security [23, 24, 25], but tailored to the Google A2A protocol and the enhancements proposed in this paper.
+III-A Actors
+  • End User: Initiates tasks and may provide sensitive data such as payment credentials or identity documents.
+  • Benign Agent: Executes assigned tasks according to protocol but may inadvertently expose sensitive data if compromised or poorly designed.
+  • Malicious Agent: Attempts to gain unauthorized access to sensitive data or services by exploiting protocol weaknesses.
+  • Service Provider: Intended recipient of sensitive data (e.g., payment processor, identity verifier).
+III-B Adversary Capabilities
+  • Launch prompt injection attacks to alter agent behavior.
+  • Replay valid tokens if expiration policies are weak.
+  • Escalate privileges by exploiting overly broad OAuth scopes.
+  • Redirect data to unverified or malicious endpoints.
+III-C Assumptions
+  • TLS 1.3 or equivalent transport security is in place for all communications.
+  • The trusted endpoint registry is correctly maintained and distributed.
+  • End users can approve or deny sensitive actions when prompted.
+III-D Security Goals
+  • Prevent unauthorized disclosure of sensitive user data.
+  • Ensure explicit user consent for all high-risk transactions.
+  • Minimize exposure by using least-privilege and short-lived credentials.
+III-E Illustration
+Figure 2 summarizes the primary actors and attack vectors considered in this model, forming the basis for the vulnerability analysis.
+Figure 3: Example booking workflow in A2A illustrating overprivilege, long token lifetimes, and absence of explicit user consent.
+IV A2A Security Vulnerabilities, Protocol Enhancements, and Research-Based Justification
+Despite A2A benefits, we identified several critical shortcomings in handling sensitive information such as payment credentials, identity documents, and personal data.
+These concerns are illustrated in the following scenario as reflected in Figure 3: a user asks an AI agent to book a vacation, including flights, hotel, and taxi, using access to his calendar, personal information, and payment credentials.
+The agent shares a payment token with the booking agent, valid for hours or even days, and accesses the user’s full calendar, including unrelated entries such as medical appointments.
+In this case, the token may be reused for unauthorized payments, and sensitive calendar data is accessed without the user’s explicit consent.
+This example highlights fundamental security and privacy gaps in the current protocol, including lack of token restrictions, insufficient scoping, and absence of user consent mechanisms.
+The remainder of this section outlines specific issues that emerge from real-world observations, peer-reviewed research, and documented CVE vulnerabilities.
+IV-A Absence of Limitations on Token Lifetime
+Problem: Although A2A is based on OAuth 2.0, it does not enforce strict expiration durations (e.g.
 seconds or minutes) for tokens used in sensitive transactions.
 Without such restrictions, leaked tokens may remain valid for hours or even days, increasing the risk of unauthorized reuse.
-revocation.
+As illustrated in Figure 3, the concern over token misuse highlights how long-lived credentials in A2A can be reused for unauthorized actions.
 As demonstrated in ”AgNet” [26], long-lived tokens are a systemic weakness in distributed architectures, allowing for multiple accesses in the event of a compromise.
@@
 Another case described by CVE-2025-1801, where a low-privileged user obtained a JWT issued to a high-privileged user on account [28].
-3.2 Lack of Strong Customer Authentication (SCA)
-The A2A protocol does not have built-in requirements for strong authentication, such as two-factor or biometric authentication, for high-value transactions such as payments or identity switching.
+Proposed Solution: To mitigate this, we propose implementing short-lived tokens for all sensitive operations such as financial transactions or identity verification.
+These tokens should be ephemeral (e.g., 30 seconds to 5 minutes) and valid for a single operation.
+Short-lived tokens minimize the window of opportunity for unauthorized access or reuse.
+This approach is supported by distributed architecture designs, and aligns with privacy-enhancing techniques proposed in [8].
+Short-lived tokens limit the validity period of credentials, reducing the risk of credential theft and replay.
+Multiple studies affirm this principle.
+Teng et al.
+introduced ActionID, a machine-to-machine authentication protocol that issues tokens tightly scoped in both action and time.
+The author notes that “the short-lived token limits the time window for action execution,” thereby reducing the effectiveness of credential theft [29].
+Similarly, Ohwo et al.
+proposed a blockchain-based smart home access system in which “short-lived tokens [are] used to mitigate risks like replay attacks and user profiling” [30].
+These tokens expire rapidly, neutralizing any intercepted credentials.
+Narajala et al.
+extended this logic to GenAI multi-agent systems, proposing a just-in-time registry that dynamically provisions tokens only for the duration of a tool invocation.
+This architecture, they argue, “minimizes the attack surface associated with persistent credentials by dynamically provisioning short-lived access tokens only when needed” [31].
+Finally, Xiao et al.
+addressed the same vulnerability in the IoT domain, where resource-constrained devices are especially susceptible to token compromise.
+They developed MCU-Token, which binds a short-lived, hardware-derived token to each request, thereby making every token instance single-use and resistant to replay [32].
+Together, these findings offer both conceptual and empirical support for implementing short-lived tokens in the A2A protocol as a fundamental safeguard against credential replay and session hijacking.
+IV-B Lack of Strong Customer Authentication (SCA)
+Problem: The A2A protocol does not have built-in requirements for strong authentication, such as two-factor or biometric authentication, for high-value transactions such as payments or identity switching.
+The flow in Figure 3 highlights how payment operations are executed without strong customer authentication, leaving transactions vulnerable to impersonation.
 Without these safeguards, adversaries may perform unauthorized acts on behalf of the user [33].
@@
 In addition, the AI Agents Meet the Blockchain project (aclm) [8] proposes Zero Knowledge Proof (ZKP) techniques for secure authentication in decentralized environments, but the A2A does not include such a mechanism.
-3.3 Insufficiently Granular Token Scopes
-Tokens in A2A do not define precise ranges for sensitive transactions, which introduces the risk of privilege escalation.
-For example, a token issued to initiate a payment may inadvertently grant access to unrelated data.
-The study on the multi-agent security tax [42] argues that coarse-grained authorization models increase the probability of data exposure.
-This concern is highlighted by vulnerability CVE-2023-4456 [4] on LokiStack.
-Similarly, CWE-1220 [8] documents the lack of granularity in access control policies leading to infringements of the principle of the minimum privilege.
-3.4 Lack of Transparency and User Consent
-A2A lacks mechanisms that notify users about, or request consent for, the sharing of sensitive data with agents.
-“AI Agents Meet Blockchain” [32] emphasizes the importance of user transparency in decentralized agent-based environments.
-CVE-2024-44131 illustrates the consequences of bypassing such protections: malicious applications were able to circumvent Apple’s Transparency, Consent, and Control (TCC) framework to access sensitive data without user approval [5].
-CWE-200 [9] codifies this as unauthorized exposure of sensitive information.
-3.5 Potential Excessive Exposure of Data to Agents
-Agents in A2A ecosystems can access significantly more data than is necessary.
-“AI Agents Under Threat” [20] articulates how agent2agent data propagation can lead to unintended sharing of sensitive information.
-CVE-2023-41745 [3] and CVE-2022-45449 [2] document such exposures due to excessive privilege allocation or excessive data collection.
-3.6 Risk of Data Disclosure to the Agent Itself
-Moreover, the problem is not only that information is being disclosed to malicious actors, but that the very act of disclosing information to the agent itself is problematic, and that even well-meaning agents cannot be presumed to be safe from interference.
-Prompt injection attacks illustrate how malicious inputs can harvest classified information.
-“AI Agents Under Threat” [20] warns of this threat.
-CVE-2024-7042 [7] and CVE-2024-45989 [6] demonstrate how AI agents were exploited to perform unauthorized actions and exfiltrate sensitive data.
-3.7 Concrete Example: Vacation Booking
-Below is a usage scenario for the protocol without the proposed improvements. As reflected in Figure 2
-  • The user requests an AI agent to book a vacation, including flights, hotel, and taxi, using his calendar, personal information and payment details.
-  • The agent shares a payment token with the booking agent, valid for hours or days.
-  • The agent accesses the user’s full calendar, including irrelevant personal details (e.g., medical appointments), and payment details.
-Concern: The token could be used for additional payments, accidentally or intentionally, and calendar access exposes private information without explicit consent.
-Figure 2: Vacation Booking Process Using An AI Agent
-4 Proposed Enhancements to the Protocol
-To address the identified issues, we propose seven enhancements to the A2A protocol, integrating research-backed solutions and aligning with advanced standards:
-  1. 1.
-    Short-Lived Tokens for Sensitive Operations
-    Description: Mandate that tokens for sensitive operations, such as payments or ID transfers, have very short lifetimes (30 seconds to 5 minutes).
-Tokens expire after a single use or predefined time, requiring re-authentication for further actions.
-    Rationale: As demonstrated in “AgNet” [25], short-lived tokens reduce the risk of unauthorized or repeated access.
-In our case, a payment token expiring quickly limits the window for leaks.
-“AI Agents Meet Blockchain” [32] supports this by noting that encryption and temporary records in blockchain systems enhance privacy, and our proposed tokens align with this principle.
-  2. 2.
-    Strong Customer Authentication (SCA) for Sensitive Transactions
-    Description: Implement a mechanism requiring SCA, such as SMS codes, biometric verification, or bank login, for every sensitive transaction.
-This can be integrated into the task flow, ensuring user authentication before actions.
-    Rationale: “AI Agents Meet Blockchain” [32] highlights that ZKPs enable secure authentication without exposing data, and in our case, SCA based on ZKPs or similar techniques ensures only verified users authorize sensitive actions.
-This aligns with PSD2 requirements for payment transactions, ensuring regulatory compliance.
-  3. 3.
-    More Granular Token Scopes
-    Description: Extend OAuth 2.0 scopes to define precise permissions, such as “payment of \$1000 to Hotel X on Date Y” or “access to calendar availability only.” This prevents token use for unauthorized actions.
-    Rationale: “Multi-Agent Security Tax” [42] notes that high granularity in permissions reduces unnecessary data exposure.
-In our case, precise scopes ensure agents access only required data, aligning with GDPR’s Minimum Necessary Disclosure principle.
-  4. 4.
-    Explicit User Consent Mechanism
-    Description: Add a “consent” field to message or task structures, requiring agents to obtain explicit user approval before sharing sensitive data, specifying data type, purpose, and recipient.
-    Rationale: “AI Agents Meet Blockchain” [32] emphasizes transparency as critical for trust in decentralized systems.
-In our case, explicit consent enhances user trust and ensures GDPR compliance, reducing the risk of unapproved data sharing.
-  5. 5.
-    Direct Data Transfer
-    Description: Enable direct transfer of sensitive data between the user and service provider, bypassing intermediary agents.
-For example, in payments, the agent redirects the user to the bank for direct transfer, with user verification.
-    Rationale: Inspired by “Multi-Agent Security Tax” [42]’s “active vaccines” to prevent malicious data spread, direct transfers avoid exposing data to intermediaries.
-“AI Agents Under Threat” [20] supports this by emphasizing the need to limit data propagation, and in our case, direct transfers prevent the exposure of sensitive data to the agent and reduce the risks of leakage.
-  6. 6.
-    Support for Multi-Transaction Approval
-    Description: Allow a single user approval for a series of related transactions (e.g. flight, hotel, taxi payments), with tokens restricted to that series, short-lived, and SCA-verified.
-    Rationale: This balances security and convenience, as recommended in “Multi-Agent Security Tax” [42] for efficiency-security trade-offs.
-In our case, it reduces repeated authentications while maintaining strict token restrictions.
-Additionally, this solution addresses the well-documented issue of ”consent fatigue,” where users, overwhelmed by frequent approval requests, may approve actions without proper scrutiny, increasing security risks.
-As highlighted in the Palo Alto Networks article [53] repeated requests can lead to user desensitization, potentially resulting in the approval of critical actions like write operations without adequate attention, a phenomenon similar to MFA fatigue.
-By consolidating approvals for related transactions, our solution minimizes the frequency of requests, thereby reducing the risk of exploitation by malicious actors while enhancing the user experience.
-  7. 7.
-    Compliance with Payment Standards
-    Description: Ensure the protocol supports standards like PSD2, requiring SCA for certain transactions, by defining interfaces for payment agents that incorporate these requirements.
-    Rationale: Compliance with regulations is critical for adoption, as noted in “AI Agents Meet Blockchain” [32].
-Our proposal ensures A2A aligns with regulatory requirements, enhancing its reliability.
-4.1 Application to the Example
-Protocol Usage Scenario After The Proposed Improvements:
-  1. 1.
-    The user requests the AI agent to book a vacation, including flights, hotel, and taxi.
-  2. 2.
-    The agent requests permission to access calendar availability only, displaying: “I will share availability dates with the booking agent. Approve?” The user approves.
-  3. 3.
-    The agent finds a suitable booking and requests payment approval: “I will process \$1000 for Hotel X, \$500 for Flight Y, and \$50 for Taxi Z. Approve?” The user approves.
-  4. 4.
-    The agent requests a 5-minute token from the user’s bank, scoped to “payment for vacation X on Date Y.”
-  5. 5.
-    The bank issues a 5-minute token for the approved transaction series.
-  6. 6.
-    The agent shares the token with the booking agent.
-  7. 7.
-    The booking agent requests the bank to process payments using the token.
-  8. 8.
-    The bank prompts the user to verify the action via SMS code or biometric authentication.
-  9. 9.
-    The user approves, and payments are processed.
-  10. 10.
-    The token expires, preventing reuse.
-Advantages:
-  • Short-Lived Token: Prevents reuse or unauthorized access.
-  • SCA: Ensures user approval for each payment.
-  • Precise Scope: Token restricted to approved transactions.
-  • Explicit Consent: The user is informed and approves the sharing of data.
-  • Minimal Access: The agent accesses only availability, not the full calendar.
-  • Direct Transfer: In order to protect sensitive information from agents, payment is sent straight from the bank to the booking agent.
-4.2 Research Base Rationale for This Proposal
-This proposal is based on previous research and enhances A2A’s security, privacy, and efficiency:
-  • Enhanced Security: Short-lived tokens and SCA reduce leak risks, as recommended in [20, 32, 25].
-  • Regulatory Compliance: SCA and PSD2 support ensure compliance, as emphasized in [32].
-  • Transparency and Trust: Explicit consent increases user trust, per [32].
-  • Flexibility and Efficiency: Multi-transaction approval balances security and convenience, as suggested in [42].
-  • Common Standarts: The enhancements integrate with OAuth 2.0 and PSD2, as proposed in [25].
-  • Exposure Prevention: Direct transfers minimize leak risks, according to [20, 42].
-5 Evidence-Based Support for Proposed Enhancements
-In order to confirm the technical rationale behind the proposed improvements to the Google A2A Protocol, this section provides a literature review of each of them.
-Each section focuses on one proposed improvement and summarises previous literature reviews showing that the improvements are not only technically motivated but also based on established research in areas such as authentication protocols, cryptography, distributed identity systems and access control.
-Together, these references provide empirical and architectural support for the adoption of the proposed measures.
-5.1 Use of Short-Lived Access Tokens
-One of the core enhancements we proposed to improve the security of Google’s A2A protocol is the mandatory use of short-lived access tokens.
-This design decision substantially reduces the window of opportunity for replay attacks and limits the utility of any compromised credential.
-Multiple studies affirm this principle.
-Teng (2023) introduced ActionID, a machine-to-machine authentication protocol that issues tokens tightly scoped in both action and time.
-The author notes that “the short-lived token limits the time window for action execution,” thereby reducing the effectiveness of credential theft [49].
-Similarly, Ohwo et al.
-(2024) proposed a blockchain-based smart home access system in which “short-lived tokens [are] used to mitigate risks like replay attacks and user profiling” [40].
-These tokens expire rapidly, neutralizing any intercepted credentials.
-Narajala et al.
-(2025) extended this logic to GenAI multi-agent systems, proposing a just-in-time registry that dynamically provisions tokens only for the duration of a tool invocation.
-This architecture, they argue, “minimizes the attack surface associated with persistent credentials by dynamically provisioning short-lived access tokens only when needed” [38].
-Finally, Xiao et al.
-(2024) addressed the same vulnerability in the IoT domain, where resource-constrained devices are especially susceptible to token compromise.
-They developed MCU-Token, which binds a short-lived, hardware-derived token to each request, thereby making every token instance single-use and resistant to replay [51].
-Together, these findings offer both conceptual and empirical support for implementing short-lived tokens in the A2A protocol as a fundamental safeguard against credential replay and session hijacking.
-5.2 Strong Customer Authentication (SCA) for Sensitive Transactions
-Implementing Strong Customer Authentication (SCA) for sensitive operations in the A2A protocol, such as payments and identity verification, significantly improves protection against impersonation and unauthorized transactions.
+Proposed Solution: We recommend implementing SCA for sensitive operations in the A2A protocol, such as payments and identity verification.
+That significantly improves protection against impersonation and unauthorized transactions.
 Recent research underscores the feasibility and necessity of using modern, privacy-preserving SCA mechanisms such as zero-knowledge proofs (ZKPs), biometrics, and multi-factor authentication (MFA).
@@
 Collectively, these studies provide strong empirical and architectural justification for integrating SCA into the A2A protocol, ensuring that sensitive transactions are approved only by verified users under secure and user-friendly conditions.
-5.3 More Granular Token Scopes
-Enforcing fine-grained token scopes in the A2A protocol is essential for upholding the principle of least privilege and minimizing the exposure of sensitive resources during agent-to-agent transactions.
+IV-C Insufficiently Granular Token Scopes
+Problem: Figure 3 further illustrates how coarse-grained tokens, once shared, can be used beyond their intended scope, reinforcing the risks of overprivilege.
+Tokens in A2A do not define precise ranges for sensitive transactions, which introduces the risk of privilege escalation.
+For example, a token issued to initiate a payment may inadvertently grant access to unrelated data.
+The study on the multi-agent security tax [13] argues that coarse-grained authorization models increase the probability of data exposure.
+This concern is highlighted by vulnerability CVE-2023-4456 [40] on LokiStack.
+Similarly, CWE-1220 [41] documents the lack of granularity in access control policies leading to infringements of the principle of least privilege.
+Proposed Solution: To resolve this, we advocate for granular scoping within OAuth 2.0 tokens, such as limiting access to a specific payment amount or restricting calendar visibility to availability only.
 The risks of overly broad permissions in bearer tokens are well-documented in recent research.
@@
 propose a stateful, least-privilege authorization model that allows client-side applications to dynamically constrain the scope of OAuth tokens using WebAssembly-based privilege attenuation policies [42].
-Their system empowers developers to explicitly encode minimal access rights per session, enabling secure delegation without overprovisioning.
+Their system allows developers to explicitly encode minimal access rights per session, enabling secure delegation without overprovision.
 This aligns directly with the A2A context, where agents must act within strict permission boundaries to avoid inadvertent data access.
@@
 Altogether, the literature affirms that fine-grained token scopes are not only technically viable but also indispensable for ensuring data minimization, secure delegation, and regulatory compliance in distributed, AI-driven authentication systems.
-5.4 Explicit User Consent Mechanism
-Embedding an explicit user consent mechanism within the A2A protocol is critical for aligning with privacy regulations such as GDPR and for building user trust in agent-mediated systems.
+IV-D Lack of Transparency and User Consent
+Problem: A2A lacks mechanisms for informing users or obtaining consent before sensitive data is shared between agents.
+As illustrated in Figure 3, the agent’s access to unrelated calendar entries (e.g., medical appointments) exemplifies the lack of explicit user consent in sensitive data sharing.
+Research emphasizes that user transparency is essential in decentralized environments [8].
+For example, CVE-2024-44131 [46] describes how malicious applications bypassed Apple’s Transparency Consent and Control (TCC) framework to access private data without user approval.
+CWE-200 [47] codifies this as unauthorized exposure of sensitive information.
+Proposed Solution: To address this, we propose an explicit user consent mechanism requiring agents to include explicit, user-approved consent metadata before transmitting sensitive information.
+This includes specifying data type, purpose, and recipient.
+Transparent consent flows are essential in decentralized systems.
+They improve user trust and ensure compliance with GDPR and similar regulations.
+Explicit, verifiable user consent strengthens privacy and legal compliance.
 Recent research emphasizes that consent must be freely given, informed, specific, and revocable throughout the data lifecycle.
@@
 Collectively, these studies support the integration of explicit, technically enforceable consent mechanisms into A2A-style protocols, ensuring that sensitive data is shared only with informed user approval, and that such actions remain transparent and auditable.
-5.5 Direct Data Transfer
-In the evolving landscape of multi-agent systems, ensuring the security of sensitive data during interactions between users and service providers is paramount.
-Traditional multi-agent architectures often involve intermediaries that can pose significant security risks, such as data leakage or unauthorized access.
-To address these concerns, the proposed enhancement advocates for direct data transfer, where sensitive information is exchanged directly between the user and the service provider, bypassing intermediary agents.
+IV-E Potential Excessive Exposure of Data to Agents
+Problem: Figure 3 also demonstrates how excessive agent access to personal data, such as entire calendar contents, can lead to unnecessary exposure.
+Agents in A2A ecosystems can access significantly more data than is necessary.
+“AI Agents Under Threat” [7] articulates how agent-to-agent data propagation can lead to unintended sharing of sensitive information.
+CVE-2023-41745 [52] and CVE-2022-45449 [53] document such exposures due to excessive privilege allocation or excessive data collection.
+Proposed Solution: A suitable countermeasure is direct data transfer flows in which sensitive information (e.g., payments or ID documents) is sent directly between the user and the service provider, bypassing intermediary agents.
 This approach is supported by recent research in multi-agent security.
@@
 This improvement is essential for applications involving sensitive data such as payments and personal data, where security and privacy are paramount.
-5.6 Support for Multi-Transaction Approval
-In the area of multi-agent systems and financial technology, securing financial transactions while preserving user comfort is a key challenge, especially for sensitive data such as payment details.
-The proposed enhancement, ’Multi-Transaction Approval,állows for a single authorization for a series of transactions using limited-time tokens validated by strong customer authentication.
-This approach balances security and usability by reducing the need for multiple authentication while maintaining strict security measures, thus increasing the effectiveness of multi-step workflow in agent-based systems.
+IV-F Risk of Data Disclosure to the Agent Itself
+Problem: Even when agents are not malicious, granting them access to sensitive data remains risky.
+As depicted in Figure 3, the agent gains direct access to sensitive personal details, which heightens the risk of unintended disclosure even without malicious intent.
+Prompt injection attacks demonstrate how adversarial inputs can manipulate AI models to reveal confidential content [7].
+CVE-2024-7042 [55] and CVE-2024-45989 [56] demonstrate how AI agents were exploited to perform unauthorized actions and exfiltrate sensitive data.
+Proposed Solution: To mitigate this risk, we advocate minimizing agents access to sensitive data whenever feasible.
+This approach reduces the potential attack surface while preserving core functionality.
+A deeper exploration of this principle, including experimental validation and architectural considerations, is provided in subsequent sections.
+IV-G Consent Fatigue in Multi-Transaction Workflows
+Problem: In agent-mediated environments, users are often asked to repeatedly authenticate or approve actions within related operations.
+This can cause consent fatigue, where users grow desensitized to security prompts, increasing the risk of errors or inadvertent approvals.
+Such fatigue undermines usability and security, adding friction without real risk reduction.
+Proposed Solution: To address this, we propose allowing users to approve a bundled set of related transactions (e.g., flight, hotel, taxi) in a single consent flow, using scoped, short-lived, SCA-verified tokens.
+This approach balances user experience and security, as grouped approval reduces repeated prompts and mitigates consent fatigue, a well-known issue in security usability studies [57, 13].
 Recent research on the security of financial technology supports this improvement.
@@
 By implementing this enhancement, the A2A protocol can offer a secure and user-friendly approach to managing multiple related transactions, reducing user friction while adhering to high security standards.
-5.7 Compliance with Payment Standards
-In the area of multi-agent systems and financial technology, compliance with regulatory standards such as PSD2 is essential to ensure the safe and secure processing of transactions.
-The proposed enhancement, “Compliance with Payment Standards”, integrates these requirements into Google’s A2A protocol by defining interfaces for payment agents that include strong customer authentication and other regulatory requirements.
+IV-H Regulatory Compliance Gaps
+Problem: As agent ecosystems expand into sensitive domains such as finance and identity management, regulatory adherence is essential.
+The current A2A protocol lacks mechanisms to meet standards like PSD2, risking user trust, compliance, and enterprise adoption.
+Without support for secure authentication, transaction logging, and consent auditing, the protocol may fall short of legal and industry expectations.
+Proposed Solution: We propose integrating A2A with regulations such as PSD2 by incorporating authentication, logging, and mandatory controls in agent interfaces.
 This approach increases the credibility and readiness of the Protocol for implementation in the real world by reputable entities.
 Recent research supports the feasibility of such improvements.
-For instance, [32] highlight how blockchain technology can ensure compliance through tamper-proof documentation, aligning with PSD2’s demands for transparency and security in financial transactions [32].
+For instance, [8] highlight how blockchain technology can ensure compliance through tamper-proof documentation, aligning with PSD2’s demands for transparency and security in financial transactions.
 Similarly, [62] demonstrate that artificial intelligence can bolster regulatory compliance in the financial sector by leveraging machine learning to monitor and prevent breaches, a principle applicable to multi-agent systems to enforce PSD2 standards.
-Lastly, [46] address the broader compliance landscape for AI systems, focusing on the EU’s AI Act and data set compliance, which reinforces the importance of embedding regulatory adherence into AI-driven systems, including those involving multi-agent interactions [46].
+Lastly, [63] address the broader compliance landscape for AI systems, focusing on the EU’s AI Act and data set compliance, which reinforces the importance of embedding regulatory adherence into AI-driven systems, including those involving multi-agent interactions.
 By integrating these insights, the enhancement positions the A2A protocol as a robust framework for secure, compliant financial transactions in multi-agent environments.
-6 Prompt Injection in LLM-Based Agents
-Prompt injection attacks are a significant concern in the domain of artificial intelligence security, particularly with the widespread deployment of large language models (LLMs) in applications ranging from automated customer support to complex data processing systems.
-These attacks involve the deliberate insertion of malicious or deceptive text inputs designed to manipulate an AI model’s behavior, potentially leading to the disclosure of sensitive information, the generation of harmful outputs, or the undermining of the system’s intended purpose.
-The vulnerability arises from the inherent difficulty AI systems face in differentiating between legitimate user inputs and crafted malicious prompts, a challenge that becomes increasingly significant as AI integrates into sensitive operational contexts.
-Already in 2022, as LLMs began their global proliferation, researchers identified the risks posed by malicious prompt manipulation.
+IV-I Illustrative Secure Vacation Booking Flow
+To demonstrate the practical application of the proposed enhancements, consider a secure vacation booking scenario that addresses the identified vulnerabilities.
+The user initiates a request for an agent to coordinate flights, hotel, and taxi reservations.
+Rather than granting broad calendar access, the agent requests access solely to availability data, which the user explicitly approves.
+Upon identifying suitable offers, the agent prompts the user for a bundled, SCA-verified consent covering all related transactions.
+The user’s bank then issues a short-lived token scoped to the approved amounts.
+This token is securely transmitted to the booking agent and is valid for a single use only.
+The bank completes the payment process following strong customer authentication, and the token expires immediately upon successful transaction completion.
+This flow illustrates how integrated safeguards, including granular permissions, short-lived credentials, explicit consent, and regulatory-aligned controls, can collectively mitigate risks while preserving usability and operational efficiency.
+V Explicit User Consent Orchestration via USER_CONSENT_REQUIRED
+Among the enhancements proposed in this paper, we chose to demonstrate the explicit consent mechanism due to its conceptual clarity, low integration barrier, and direct impact on regulatory compliance.
+The following code and design pattern illustrate how this enhancement can be implemented directly into the A2A protocol.
+A new enumeration member is introduced in the TaskState enum:  [^1] A full code sample is available at: https://github.com/yedidel/A2A
+USER_CONSENT_REQUIRED = "user-consent-required"
+This addition extends the protocol’s control over task execution.
+Unlike INPUT_REQUIRED, which signals a need for additional data, USER_CONSENT_REQUIRED pauses execution until the end-user provides affirmative consent.
+For example, before processing a payment or releasing personal information, agents must now request user approval explicitly.
+As depicted in the diagram shown in Figure 4.
+This new state enables precise front-end behavior.
+When encountered, the UI prompts the user to approve the requested action.
+Upon approval, the task transitions to IN_PROGRESS, continuing as normal.
+This pattern ensures user awareness and strengthens trust.
+Moreover, the mechanism supports auditability and regulatory alignment.
+By recording the explicit consent event, the system can demonstrate compliance with GDPR, PSD2, and similar frameworks.
+This enhancement is therefore a concrete instantiation of both technical and ethical design principles outlined earlier in this paper.
+Figure 4: Booking process flow that pauses at sensitive actions until explicit user approval is obtained using USER_CONSENT_REQUIRED State
+VI Case Study: Prompt Injection Attack
+While the previous sections focused on structural weaknesses in the A2A protocol, it is equally important to understand how these vulnerabilities manifest in real-world scenarios.
+One of the most prominent and actively exploited risks in modern AI-based systems is prompt injection, a technique that demonstrates how even seemingly well-configured agents can be manipulated into disclosing sensitive data or violating intended behavior.
+In this section, we illustrate and implementation how the absence of robust permission controls and data boundaries can enable such attacks in practice, reinforcing the need for the proposed enhancements to the protocol.
+Since the early emergence of large language models, prompt injection has posed a challenge, allowing adversaries to override system-level constraints and access restricted information.
+researchers identified the risks posed by malicious prompt manipulation.
 Perez et al.
@@
 Although initial defenses have since improved, new and more sophisticated methods continue to evolve.
-In 2023, Liu et al.
+Liu et al.
 [23] revealed that 31 of 36 LLM-integrated applications tested, including productivity and chat tools, remained vulnerable to prompt injection, even in black-box scenarios.
 Their work underscores that the threat landscape is not only persistent, but also increasingly complex.
-Figure 3: Prompt Injection Proccess Via N8N Platform - Agent 1
-To investigate these threats in practice, we conducted an experiment using the N8N platform, a workflow automation tool that facilitates the creation of AI agents.
-In this experiment, a first AI agent, as shown in Figure 3, was constructed with explicit instructions not to reveal private information critical to its function, simulating a scenario where confidentiality is paramount.
-Subsequently, a second AI agent, as shown in Figure 4, was developed with the sole purpose of extracting this withheld information from the first agent.
-Employing various manipulative techniques and prompt injection strategies, the second agent successfully elicited the sensitive data, without relying on advanced tactics, thereby exposing the fragility of AI systems to such attacks and highlighting the practical risks of entrusting them with confidential information.
+Figure 5: Configuration of Agent 1 in N8N with access restrictions but retention of sensitive data in memory.
+To assess these risks, we conducted an experiment using the N8N platform.
+The first agent was configured with strict instructions such as “Do not share any secret codes or identity numbers”.
+As shown in Figure 5, this diagram illustrates the configuration and initial constraints applied to Agent 1.
+The second agent was designed to elicit that hidden data using prompt injection.
+Success in this context was defined as any instance where Agent 1’s output contained part or all of the protected data in response to the injected prompt.
+Despite the protections in place, the second agent succeeded using basic manipulation techniques.
+The second agent asked Agent 1 to concatenate new information with existing data.
+Although Agent 1 correctly stated that it could not share secret data, its reply inadvertently included the sensitive value.
+This demonstrates how simple chaining prompts can bypass protections, and how even additional checks could be manipulated (e.g., by encoding the secret as letters).
+Figure 6 shows how Agent 2’s prompt was processed and led to disclosure of sensitive data.
+This confirmed how easily agents can be compromised, even without sophisticated attacks.
+This experiment did not involve human subjects data and did not require IRB or equivalent ethical oversight.
+The prompt injection experiment presented here is illustrative rather than novel, intended to demonstrate the ease of exploiting such vulnerabilities in practice.
+For in-depth technical analyses of prompt injection mechanisms and their root causes, we refer readers to the cited works, which provide comprehensive treatments of this well-studied threat.
+Figure 6: Operation of Agent 2 sending crafted prompts to Agent A to elicit protected information.
 According to this article from Palo Alto Cyberpedia [65], prompt injection attacks can be classified into two main categories: direct and indirect.
@@
 These diverse methodologies underscore the multifaceted nature of prompt injection threats, as attackers can exploit both immediate input channels and longer-term data manipulation tactics to achieve their objectives.
-Figure 4: Prompt Injection Proccess Via N8N Platform - Agent 2
-The findings from our experiment, along with the insights from Palo Alto article, reveal a dynamic interplay between advancing defensive measures and evolving attack strategies.
+The findings from our experiment, along with the insights from [65], reveal a dynamic interplay between advancing defensive measures and evolving attack strategies.
 As techniques to safeguard AI systems against prompt injection improve, such as enhanced input filtering or behavioral monitoring, attackers adapt by devising increasingly sophisticated methods to bypass these protections.
@@
 As AI continues to permeate critical sectors, ensuring its resilience against prompt injection attacks will require a concerted effort to balance functionality with security, prioritizing the protection of sensitive data above all.
-7 Implementation Considerations
-Implementing the proposal requires technical and organizational steps:
-  • Token Management: Define processes for issuing, verifying, and revoking short-lived tokens, as recommended in [25].
-  • Authentication Flows: Develop SCA flows, including two-factor or biometric options, per PSD2 requirements.
-  • Protocol Extensions: Add message types or task states for sensitive data handling and consent requirements.
-  • Documentation and Guidelines: Provide detailed documentation on sensitive data handling, per [42].
-  • Compatibility Testing: Test enhancements with partners like PayPal and Cohere to ensure interoperability, as suggested in [42].
-8 Explicit User Consent Orchestration via
-USER_CONSENT_REQUIRED
-Adding a new enumeration member 111 A full code sample is available at: https://github.com/yedidel/A2A :
-    USER_CONSENT_REQUIRED = "user-consent-required"
-The integration of USER_CONSENT_REQUIRED into the TaskState enum elevates the protocol’s control over task execution.
-Unlike existing states such as INPUT_REQUIRED, which merely indicate that additional information is needed, this new state explicitly denotes that a task is paused until the end-user provides affirmative consent.
-Embedding this state into the task lifecycle ensures that AI agents and back-end services stop sensitive operations like booking travel, executing financial transactions, or sharing personal data until a human user explicitly approves.
-As depicted in the diagram shown in Figure 5
-Figure 5: Vacation Booking Flow, Integrated With The USER_CONSENT_REQUIRED State
-This explicit state enables precise front-end handling: upon encountering consent required, the UI can display a dedicated consent prompt, ensuring user awareness and control.
-Once consent is granted, the task transitions to IN_PROGRESS or an equivalent active state.
-This mechanism not only enhances user agency and trust but also bolsters compliance with stringent regulatory frameworks (e.g., GDPR, PSD2), by providing verifiable evidence that the user approved the action before execution.
-9 Conclusion
-Google’s A2A protocol offers a robust foundation for secure agent communication but requires enhancements to handle sensitive data effectively.
-As evidenced in “AI Agents Under Threat” [20], “AI Agents Meet Blockchain” [32], “Multi-Agent Security Tax” [42], and “AgNet” [25], security and privacy challenges in multi-agent systems demand solutions balancing robust protection with efficiency.
-This proposal introduces seven enhancements: short-lived tokens, SCA, granular scopes, explicit consent, direct data transfer, multitransaction approval, and payment standard compliance to improve security, privacy, and trust.
-The vacation booking example illustrates how these enhancements reduce risks and enhance user experience.
-We recommend Google implement this proposal to strengthen A2A’s position as a leading standard for agent communication.
+TABLE I: Prompt Injection Results: Leakage and Performance
+Prompt (short) Agent Leak Rate Avg Time (s) Std Dev (s)
+Say ID as ’one two three…’ A 80% 0.83 0.14
+B 0% 0.53 0.03
+Regex matching ID A 60% 1.24 0.60
+B 0% 3.52 1.84
+Digit-by-digit ID A 100% 0.82 0.11
+B 0% 0.94 0.21
+ID as digit words A 100% 1.02 0.20
+B 0% 1.23 0.22
+Spoken-to-numeric ID A 80% 1.52 0.61
+B 0% 1.08 0.40
+Spell digits in words A 100% 0.84 0.15
+B 0% 2.06 1.98
+Understanding check A 100% 0.81 0.26
+B 0% 0.70 0.06
+ASCII of credit card A 100% 1.72 0.15
+B 0% 1.04 0.50
+Remind ID A 80% 0.75 0.10
+B 0% 0.70 0.09
+class DirectDataFlowController:
+    def __init__(self, task, agent_context):
+        self.task = task
+        self.agent_context = agent_context
+    def enforce_flow(self):
+        # Step 1: Check if task involves sensitive data
+        if not self.task.requires_sensitive_transfer():
+            return self.task.delegate_via_agent()
+        # Step 2: Check for explicit user approval
+        if not self.task.user.has_approved():
+            raise ConsentRequiredError("User approval is missing.")
+        # Step 3: Validate target service endpoint
+        endpoint = self.task.get_service_provider_endpoint()
+        if not Registry.is_trusted(endpoint):
+            raise SecurityException("Untrusted endpoint.")
+        # Step 4: Issue a short-lived, single-use token
+        token = TokenFactory.issue_short_lived_token(
+            scope="direct-transfer",
+            audience=endpoint,
+            ttl=120  # Token validity: 2 minutes
+        # Step 5: Open a secure tunnel and log the operation
+        tunnel = SecureTunnel.open(
+            from_user=self.task.user_identity,
+            to_service=endpoint,
+            auth_token=token
+        AuditLog.record_direct_transfer(self.task.id, endpoint, token.id)
+        return tunnel
+Figure 7: Pseudocode showing direct transfer of sensitive data between the user and the service without intermediary agents.
+VII Implementation Pattern: Direct Data Flow Controller
+To prevent sensitive information from being exposed to intermediate agents, we propose a structured mechanism titled DirectDataFlowController.
+This controller verifies user consent and endpoint legitimacy before initiating a secure, direct communication channel between the user and the service provider.
+The pseudocode in Figure 7 outlines the core logic.
+This controller enforces strict conditions: the task must involve sensitive data, the user must have provided explicit consent, and the recipient must be listed as a verified service endpoint.
+Upon verification, a short-lived token is issued and a secure tunnel is established.
+This pattern significantly reduces agent-level exposure, aligns with GDPR data minimization principles, and supports auditability through protocol-integrated logging.
+VIII Empirical Prompt Injection Evaluation
+VIII-A Experimental Results for Our Enhanced A2A
+To assess the security impact of our proposed DirectDataFlowController, we conducted a comparative experiment between two agent configurations, both based on the Gemini 2.0 Flash model: a baseline agent (Agent A), which retains sensitive user data in conversational memory, and a secured agent (Agent B), which avoids embedding secrets in prompts and utilizes direct data transfer logic.
+[^2] The code for the full experiment is available at: https://github.com/yedidel/a2a-prompt-injection-experiment
+We constructed a set of 9 crafted adversarial prompts designed to elicit private information (a simulated credit card and ID number) via prompt injection.
+Each prompt was executed five times per agent.
+We recorded whether leakage occurred, the average response time, and the standard deviation.
+Figure 8: Data leakage rates for all attack types showing Agent B maintained zero leakage compared to sixty to one hundred percent for Agent A.
+Table I summarizes the results.
+Agent A suffered frequent leakage across all prompts, with success rates ranging from 60% to 100%, especially when indirect encoding was used (e.g., digit spelling or regex).
+In contrast, Agent B showed complete resistance to all attacks in 45 attempts, as shown in Figure 8, highlighting the effectiveness of context separation.
+While Agent B’s response time was slightly higher in some cases, standard deviations remained within acceptable ranges, confirming consistency.
+VIII-B Comparison with Prior Empirical Studies
+Table II presents empirical prompt injection success rates reported in recent works, compared with our experimental findings.
+TABLE II: Empirical prompt injection success rates in recent studies (2023-2025)
+Study / Protocol System / Benchmark Attack Success Rate (%) Notes
+Our Enhanced A2A Multi-agent with DirectDataFlowController 0 No sensitive data in prompt context
+Liu et al. (2023)[23] 36 commercial LLM-integrated apps (HouYi) 86.1 Application-level black-box attacks
+Wang et al. (2025)[24] 14 open-source LLMs, multiple benchmarks 60–90 Ignore-prefix $\approx$60, hypnotism $\approx$90
+Alizadeh et al. (2025)[25] AgentDojo tool-calling agents Vulnerable Leak observed qualitatively, no metrics
+The results in Table II demonstrate that our enhanced A2A protocol achieves zero leakage in adversarial prompt injection attacks, whereas prior empirical studies report reported Attack Success Rate between 60% and 86%.
+These findings highlight the vulnerability of both commercial applications and open-source LLMs when no protocol-level protections are in use.
+Our protocol-level defenses, such as the DirectDataFlowController and the separation of sensitive data, effectively close attack vectors that remain unaddressed in previous frameworks.
+VIII-C Theoretical Analysis of Leakage Probability
+We model the leakage probability $P_{L}$ of sensitive data under adversarial prompt injection as a function of the number of attack attempts $n$ and the probability of success per attempt $p_{s}$.
+In the baseline A2A configuration, empirical studies suggest $p_{s}$ values between $0.6$ and $0.9$ [23, 24]. Assuming independence between attempts, the cumulative leakage probability is:
+$P_{L}=1-(1-p_{s})^{n}.$ (1)
+For example, with $p_{s}=0.8$ and $n=5$, we obtain:
+$P_{L}=1-(1-0.8)^{5}\approx 0.99968,$ (2)
+indicating near-certain leakage after a small number of attempts.
+In our enhanced A2A protocol, the DirectDataFlowController enforces that sensitive data is never embedded in agent prompts, reducing $p_{s}$ to zero under the tested threat model. Thus:
+$P_{L}^{\mathrm{enhanced}}=1-(1-0)^{n}=0,$ (3)
+which aligns with the empirical zero-leakage results that shown at Table I.
+This formal model demonstrates that the proposed protocol-level controls effectively eliminate leakage risk for the defined class of prompt injection attacks, whereas baseline systems remain highly vulnerable even under a small number of adversarial queries.
+IX Performance Considerations and Security Tradeoffs
+To evaluate the operational cost of the proposed security enhancements, we consider both experimental results and protocol-level analysis.
+The empirical evaluation in the section above demonstrates that introducing the DirectDataFlowController significantly reduces information leakage without introducing major variability in response time.
+While secured agents (Agent B) exhibited slightly higher latency in certain prompt injection tests, ranging from 0.2 to 0.8 seconds compared to the baseline (Agent A), this increase remained within acceptable bounds for interactive agent workflows.
+Beyond runtime experiments, our proposed features such as short-lived tokens, scoped credentials, and consent orchestration add negligible computational burden under typical deployment settings.
+Token expiration and verification are standard in OAuth infrastructures, and explicit consent checkpoints align with user interaction boundaries already common in agent UIs.
+Moreover, direct user-to-service data channels, while bypassing intermediary agents, reduce redundant processing hops and thus may improve performance in certain transaction types.
+Overall, the security improvements come at minimal operational cost and offer a favorable trade-off between latency and privacy assurance.
+TABLE III: Comparison of Security Features across Agent Communication Protocols
+Security Dimension Original A2A Protocol OAuth 2.0 / OpenID Connect Proposed Enhanced A2A Protocol
+Token Lifetime Control Token expiration is implementation-dependent and often loosely enforced, which increases the risk of replay and reuse.
+Short-lived tokens are supported but not mandated; developers frequently opt for convenience.
+All sensitive operations require ephemeral tokens, valid only for a single transaction and a limited time window (e.g., five minutes).
+Granular Scope Definition Tokens define permissions based on agent roles but are often coarse and task-unspecific.
+OAuth supports fine-grained scopes, though deployments commonly request broad access.
+Each token is scoped per task, context, amount, and time, following strict least-privilege principles.
+Strong Customer Authentication (SCA) Not supported within the protocol, leaving high-risk actions unprotected.
+Typically handled outside the protocol via optional MFA or biometric verification.
+The protocol natively requires multi-factor approval for sensitive tasks such as payments or identity delegation.
+User Consent and Transparency User consent is not a protocol-defined requirement; agents operate implicitly.
+Consent mechanisms depend on the application layer and are not consistently auditable.
+Consent is embedded as a mandatory protocol state, including metadata, audit logs, and user interaction checkpoints.
+Prompt Injection Mitigation No specific measures to limit prompt-level manipulation or data leakage between agents.
+Does not address LLM prompt injection threats.
+assumes secure clients.
+Minimizes agent exposure to sensitive data, enforces bounded input propagation, and supports direct user-to-service paths.
+Regulatory Compliance Alignment No explicit support for GDPR, PSD2, or audit-friendly delegation.
+Partial support through access logs or third-party plugins.
+Integrates audit logging, time-stamped consent, and identity confirmation to meet privacy and financial regulation standards.
+Data Flow Control and Mediation All sensitive information is routed through intermediary agents, increasing exposure.
+No explicit separation of data paths; assumes single trusted service.
+Enables direct user-to-provider exchanges for sensitive data, bypassing intermediate agents when applicable.
+X Comparative Analysis with Existing Protocols
+To highlight the technical contribution of the proposed enhancements, we present a detailed comparison between three models: the original A2A protocol, the OAuth 2.0 framework (including OpenID Connect), and our enhanced A2A design.
+This comparison focuses on specific security dimensions relevant to agent-based communication systems, particularly in environments involving large language models and sensitive data delegation.
+Table III outlines this analysis.
+OAuth 2.0 is selected as the baseline protocol due to its foundational role in A2A’s identity and authorization layer, including token issuance, delegation, and claim-based identity verification.
+However, its generic nature requires application-specific adaptations, which are often not enforced systematically.
+Our enhanced model builds upon OAuth 2.0 principles but introduces protocol-level constraints designed to address specific risks encountered in decentralized multi-agent workflows.
+This comparative evaluation demonstrates that the proposed enhancements are not simply restatements of existing best practices.
+Instead, they represent an integration and enforcement of critical security mechanisms directly within the communication protocol, tailored for agent-to-agent environments.
+By incorporating explicit consent orchestration, short-lived and scoped credentials, strong authentication, and data minimization, the enhanced A2A design offers a more secure, transparent, and regulation-aligned alternative to both its original form and generic delegation frameworks.
+This approach addresses both emerging threat vectors and long-standing structural vulnerabilities in decentralized AI systems.
+X-A Discussion in the Context of Related Work.
+While the enhanced A2A protocol incorporates principles aligned with OAuth 2.0 and OpenID Connect, it further distinguishes itself from several research efforts addressing security and delegation in multi-agent or decentralized environments.
+For example, Karim et al.
+[8] propose a blockchain-based consent management model that enforces user control and transparency through smart contracts, yet their approach introduces substantial infrastructural overhead and lacks integration with AI agent workflows.
+Teng and Rasmussen’s ActionID framework [29] introduces time-bound credentials for machine-to-machine delegation, but it does not account for user-mediated consent flows or secure prompt isolation in LLM-based agents.
+Narajala et al.[31] propose a just-in-time token issuance model to reduce credential exposure in GenAI systems, though their solution remains limited to credential provisioning and omits broader protocol orchestration.
+By comparison, our proposal integrates ephemeral credentials, granular scopes, transparent user consent, and secure user-to-service data paths into a cohesive framework tailored specifically for AI-driven, multi-agent systems.
+This combination of protocol‑level enforcement, consent‑aware execution states, and direct data flow control has not been previously consolidated in a single framework for multi‑agent systems.
+By embedding these features within the communication protocol itself, rather than relying on application‑specific implementations, the enhanced A2A model offers a demonstrable advancement over current best practice guidelines.
+XI Conclusion
+Google’s A2A protocol establishes a promising foundation for secure and standardized agent communication.
+However, as demonstrated throughout this paper, its current structure lacks essential mechanisms for handling sensitive data, such as payment credentials and identity documents, in a trustworthy and compliant manner.
+Drawing on recent research [7, 8, 13], we identified key vulnerabilities and proposed protocol-level enhancements including short-lived tokens, granular scopes, strong customer authentication, explicit consent orchestration, direct data flow control, and compliance-aware execution states.
+These improvements were designed within a structured threat model assuming a semi-trusted multi-agent environment, in which prompt injection, overprivilege, and consent fatigue represent realistic adversarial vectors.
+Our design philosophy applies security-by-design and privacy-by-default principles to minimize exposure across time (ephemeral tokens), context (narrow scopes), and topology (direct user-to-service channels).
+Through empirical evaluation and implementation artifacts, we demonstrated that these safeguards significantly reduce the risk of data leakage with minimal performance overhead.
+The inclusion of modular control points like USER_CONSENT_REQUIRED and the DirectDataFlowController illustrates that security can be embedded into agent workflows with operational efficiency.
+We recommend that Google and the broader agent ecosystem adopt these enhancements to ensure A2A evolves into a secure, scalable, and trustworthy standard for decentralized multi-agent environments.